# Detection of Gravitational Anomaly at Low Acceleration from a Highest-quality Sample of 36 Wide Binaries with Accurate 3D Velocities

### v1

@@ E.5 Speckle Holography
Gaia DR3 5181911238426924288 2024.8619 0.42 0.29
Gaia DR3 53720557585606912 2024.8621 1.14 0.73
Gaia DR3 53720561882110336 2024.8621 1.04 0.62
Gaia DR3 5737419559713009024 2024.8625 2.01 1.69
Gaia DR3 5737419559713009152 2024.8625 1.11 0.69
Gaia DR3 5741344919302959104 2024.8679 0.24 0.18
Gaia DR3 5741344919302959360 2025.2123 0.29 0.23
Gaia DR3 5741345125461458944 2024.8679 0.25 0.19
Gaia DR3 5741345125461459200 2025.2123 0.26 0.20
Gaia DR3 5755402484702384128 2024.8625 0.42 0.29
Gaia DR3 5755402720924702720 2024.8625 0.76 0.43
Gaia DR3 5760475695776013568 2024.8625 1.44 1.02
Gaia DR3 5760475695777562880 2024.8625 1.37 1.00
Gaia DR3 579529081235551360 2024.8679 0.19 0.12
Gaia DR3 579529081235551488 2024.8679 0.18 0.12
Gaia DR3 594453852070106624 2024.8679 0.19 0.11
Gaia DR3 594453852070765696 2025.2176 0.25 0.18
Gaia DR3 601946577137602816 2024.8625 2.29 1.90
Gaia DR3 601946955094723968 2024.8625 0.77 0.43
Gaia DR3 619889232512875648 2024.8679 0.25 0.18
Gaia DR3 619890709981625344 2024.8679 0.28 0.20
Gaia DR3 6321593106912540288 2025.4340 0.53 0.36
Gaia DR3 6321593111208064384 2025.4340 0.42 0.30
Gaia DR3 6321929729270985472 2025.4339 0.24 0.19
Gaia DR3 6321929733564266752 2025.4339 0.18 0.12
Gaia DR3 633627316769035264 2024.8679 0.40 0.27
Gaia DR3 633627321064388096 2024.8679 0.33 0.23
Gaia DR3 639517366199833344 2024.8679 0.30 0.21
Gaia DR3 639517366199833472 2024.8679 0.27 0.19
Gaia DR3 644549800855341184 2024.8680 0.25 0.18
Too weak Gaia DR3 644549869574817792 2024.8680 >2.>2. >2.>2.
Gaia DR3 644936691509435520 2024.8680 0.33 0.23
Gaia DR3 644936691509435648 2025.2177 0.37 0.26
Gaia DR3 648508863053616384 2025.2177 0.28 0.20
Gaia DR3 648508867348963840 2024.8680 0.60 0.36
@@
Gaia DR3 656657622902671488 2024.8652 0.66 0.37
Gaia DR3 661127824863058176 2024.1279 0.24 0.18
Gaia DR3 661128752575993600 2024.8652 0.28 0.20
@@
Gaia DR3 668947001804697472 2024.8652 0.45 0.30
Gaia DR3 675639449909479040 2024.8652 1.28 0.83
Gaia DR3 675639449911182336 2024.8652 1.30 0.86
Gaia DR3 680661782802091392 2025.2175 0.58 0.34
Gaia DR3 680662573076074752 2024.8652 0.32 0.22
Gaia DR3 6840365615137220608 2024.6838 0.29 0.22
Gaia DR3 6840365718216434688 2024.6838 0.22 0.16
Gaia DR3 6903444700303516160 2024.6782 0.46 0.32
Gaia DR3 6903444700303516800 2024.6782 0.95 0.50
Gaia DR3 6906039925703317248 2024.6781 0.29 0.21
Gaia DR3 6906039925703317376 2024.6781 0.29 0.21
Gaia DR3 6909250362214709248 2024.6839 0.38 0.27
Gaia DR3 6909250362214915072 2024.6839 0.35 0.25
Gaia DR3 6912178464759019520 2024.6839 0.38 0.27
Gaia DR3 6912178464759020544 2024.6839 0.45 0.32
Gaia DR3 695440421670598400 2024.8680 0.31 0.22
Gaia DR3 695440421670598656 2024.8680 0.32 0.21
Gaia DR3 710817744619067520 2024.8652 1.50 1.24
Gaia DR3 710817744619067648 2024.8652 0.40 0.26
Gaia DR3 758958211973432704 2025.2179 0.26 0.19
Gaia DR3 758958929232265472 2025.2179 0.43 0.29
Gaia DR3 77161217776670208 2024.8617 0.30 0.21
Gaia DR3 77161222072044288 2024.8617 0.28 0.19
Gaia DR3 787833483266126080 2025.2123 0.34 0.25
Gaia DR3 787833551986183168 2025.2123 0.31 0.22
Gaia DR3 796311542548505984 2024.8680 0.27 0.19
Gaia DR3 796311954865365888 2024.1279 1.19 0.89
Gaia DR3 803485581043936768 2025.2123 0.41 0.28
Gaia DR3 803485615403673728 2025.2123 0.30 0.22
Gaia DR3 817937454502566528 2024.8680 0.41 0.28
Gaia DR3 817937458796755200 2024.8680 0.53 0.34
Gaia DR3 820692212165053056 2024.8680 0.19 0.12
Gaia DR3 820692216460834944 2025.2177 0.20 0.14
Gaia DR3 853328573431352576 2025.2178 0.39 0.28
Gaia DR3 853328573431352704 2025.2178 0.32 0.24
Gaia DR3 87111458705996800 2024.8617 0.97 0.58
@@
Gaia DR3 883478861596266880 2024.8623 0.75 0.41
Gaia DR3 883478865892528128 2024.8623 1.08 0.69
Gaia DR3 898728439334815616 2024.8650 0.58 0.35
Gaia DR3 898728439334815744 2024.8650 0.39 0.26
Gaia DR3 899627393169280384 2024.8651 0.42 0.26
Gaia DR3 899627393169280512 2024.8651 0.44 0.29
Gaia DR3 904299046277136768 2024.8652 1.44 0.99
Gaia DR3 904299046277137024 2024.8652 0.54 0.33
Gaia DR3 914241517609344128 2024.8652 0.26 0.17
Gaia DR3 914244399532441472 2024.8652 0.25 0.17
Gaia DR3 916482150508905856 2024.8652 0.19 0.12
Gaia DR3 917982605923642752 2024.8652 0.24 0.16
Gaia DR3 922595430863222144 2024.1279 0.87 0.59
Gaia DR3 922601585552127104 2024.1279 0.28 0.21
Gaia DR3 930712128780483840 2024.8652 1.68 1.36
Gaia DR3 930712133074729856 2024.8652 1.63 1.25
Gaia DR3 937667952870488704 2024.8623 0.23 0.16
Gaia DR3 937667952871051776 2025.2174 0.23 0.17
Gaia DR3 951148377744097024 2024.8651 0.38 0.26
Gaia DR3 951148446463574016 2024.8651 0.35 0.24
Gaia DR3 951621305182882432 2024.8650 2.64 2.10
Gaia DR3 951621305182882944 2024.8650 1.87 1.37
Gaia DR3 953178312430077312 2024.8651 1.54 1.16
Gaia DR3 953178312430645376 2024.8651 0.55 0.34
@@
Gaia DR3 982674434128011904 2024.8651 0.54 0.35
Gaia DR3 98692339803443328 2024.6869 0.23 0.15
Gaia DR3 98692614681349248 2024.6869 0.14 0.11
Gaia DR3 990524191236435200 2024.8651 0.48 0.32
Gaia DR3 990524191236435456 2024.8651 0.39 0.27
Gaia DR3 992789150829982848 2024.1278 0.28 0.21
Gaia DR3 992789872384490496 2024.8650 0.29 0.21

### v2

@@ E.5 Speckle Holography
Gaia DR3 5181911238426924288 2024.8619 0.210 0.145
Gaia DR3 53720557585606912 2024.8621 0.570 0.365
Gaia DR3 53720561882110336 2024.8621 0.520 0.310
Gaia DR3 5737419559713009024 2024.8625 1.005 0.845
Gaia DR3 5737419559713009152 2024.8625 0.555 0.345
Gaia DR3 5741344919302959104 2024.8679 0.120 0.090
Gaia DR3 5741344919302959360 2025.2123 0.145 0.115
Gaia DR3 5741345125461458944 2024.8679 0.125 0.095
Gaia DR3 5741345125461459200 2025.2123 0.130 0.100
Gaia DR3 5755402484702384128 2024.8625 0.210 0.145
Gaia DR3 5755402720924702720 2024.8625 0.380 0.215
Gaia DR3 5760475695776013568 2024.8625 0.720 0.510
Gaia DR3 5760475695777562880 2024.8625 0.685 0.500
Gaia DR3 579529081235551360 2024.8679 0.095 0.060
Gaia DR3 579529081235551488 2024.8679 0.090 0.060
Gaia DR3 594453852070106624 2024.8679 0.095 0.055
Gaia DR3 594453852070765696 2025.2176 0.125 0.090
Gaia DR3 601946577137602816 2024.8625 1.145 0.950
Gaia DR3 601946955094723968 2024.8625 0.385 0.215
Gaia DR3 619889232512875648 2024.8679 0.125 0.090
Gaia DR3 619890709981625344 2024.8679 0.140 0.100
Gaia DR3 6321593106912540288 2025.434 0.265 0.180
Gaia DR3 6321593111208064384 2025.434 0.210 0.150
Gaia DR3 6321929729270985472 2025.4339 0.120 0.095
Gaia DR3 6321929733564266752 2025.4339 0.090 0.060
Gaia DR3 633627316769035264 2024.8679 0.200 0.135
Gaia DR3 633627321064388096 2024.8679 0.165 0.115
Gaia DR3 639517366199833344 2024.8679 0.150 0.105
Gaia DR3 639517366199833472 2024.8679 0.135 0.095
Gaia DR3 644549800855341184 2024.868 0.125 0.090
Too weak Gaia DR3 644549869574817792 2024.868 >2.>2. >2.>2.
Gaia DR3 644936691509435520 2024.868 0.165 0.115
Gaia DR3 644936691509435648 2025.2177 0.185 0.130
Gaia DR3 648508863053616384 2025.2177 0.140 0.100
Gaia DR3 648508867348963840 2024.868 0.300 0.180
@@
Gaia DR3 656657622902671488 2024.8652 0.330 0.185
Gaia DR3 661127824863058176 2024.1279 0.120 0.090
Gaia DR3 661128752575993600 2024.8652 0.140 0.100
@@
Gaia DR3 668947001804697472 2024.8652 0.225 0.150
Gaia DR3 675639449909479040 2024.8652 0.640 0.415
Gaia DR3 675639449911182336 2024.8652 0.650 0.430
Gaia DR3 680661782802091392 2025.2175 0.290 0.170
Gaia DR3 680662573076074752 2024.8652 0.160 0.110
Gaia DR3 6840365615137220608 2024.6838 0.145 0.110
Gaia DR3 6840365718216434688 2024.6838 0.110 0.080
Gaia DR3 6903444700303516160 2024.6782 0.230 0.160
Gaia DR3 6903444700303516800 2024.6782 0.475 0.250
Gaia DR3 6906039925703317248 2024.6781 0.145 0.105
Gaia DR3 6906039925703317376 2024.6781 0.145 0.105
Gaia DR3 6909250362214709248 2024.6839 0.190 0.135
Gaia DR3 6909250362214915072 2024.6839 0.175 0.125
Gaia DR3 6912178464759019520 2024.6839 0.190 0.135
Gaia DR3 6912178464759020544 2024.6839 0.225 0.160
Gaia DR3 695440421670598400 2024.868 0.155 0.110
Gaia DR3 695440421670598656 2024.868 0.160 0.105
Gaia DR3 710817744619067520 2024.8652 0.750 0.620
Gaia DR3 710817744619067648 2024.8652 0.200 0.130
Gaia DR3 758958211973432704 2025.2179 0.130 0.095
Gaia DR3 758958929232265472 2025.2179 0.215 0.145
Gaia DR3 77161217776670208 2024.8617 0.150 0.105
Gaia DR3 77161222072044288 2024.8617 0.140 0.095
Gaia DR3 787833483266126080 2025.2123 0.170 0.125
Gaia DR3 787833551986183168 2025.2123 0.155 0.110
Gaia DR3 796311542548505984 2024.868 0.135 0.095
Gaia DR3 796311954865365888 2024.1279 0.595 0.445
Gaia DR3 803485581043936768 2025.2123 0.205 0.140
Gaia DR3 803485615403673728 2025.2123 0.150 0.110
Gaia DR3 817937454502566528 2024.868 0.205 0.140
Gaia DR3 817937458796755200 2024.868 0.265 0.170
Gaia DR3 820692212165053056 2024.868 0.095 0.060
Gaia DR3 820692216460834944 2025.2177 0.100 0.070
Gaia DR3 853328573431352576 2025.2178 0.195 0.140
Gaia DR3 853328573431352704 2025.2178 0.160 0.120
Gaia DR3 87111458705996800 2024.8617 0.485 0.290
@@
Gaia DR3 883478861596266880 2024.8623 0.375 0.205
Gaia DR3 883478865892528128 2024.8623 0.540 0.345
Gaia DR3 898728439334815616 2024.865 0.290 0.175
Gaia DR3 898728439334815744 2024.865 0.195 0.130
Gaia DR3 899627393169280384 2024.8651 0.210 0.130
Gaia DR3 899627393169280512 2024.8651 0.220 0.145
Gaia DR3 904299046277136768 2024.8652 0.720 0.495
Gaia DR3 904299046277137024 2024.8652 0.270 0.165
Gaia DR3 914241517609344128 2024.8652 0.130 0.085
Gaia DR3 914244399532441472 2024.8652 0.125 0.085
Gaia DR3 916482150508905856 2024.8652 0.095 0.060
Gaia DR3 917982605923642752 2024.8652 0.120 0.080
Gaia DR3 922595430863222144 2024.1279 0.435 0.295
Gaia DR3 922601585552127104 2024.1279 0.140 0.105
Gaia DR3 930712128780483840 2024.8652 0.840 0.680
Gaia DR3 930712133074729856 2024.8652 0.815 0.625
Gaia DR3 937667952870488704 2024.8623 0.115 0.080
Gaia DR3 937667952871051776 2025.2174 0.115 0.085
Gaia DR3 951148377744097024 2024.8651 0.190 0.130
Gaia DR3 951148446463574016 2024.8651 0.175 0.120
Gaia DR3 951621305182882432 2024.865 1.320 1.050
Gaia DR3 951621305182882944 2024.865 0.935 0.685
Gaia DR3 953178312430077312 2024.8651 0.770 0.580
Gaia DR3 953178312430645376 2024.8651 0.275 0.170
@@
Gaia DR3 982674434128011904 2024.8651 0.270 0.175
Gaia DR3 98692339803443328 2024.6869 0.115 0.075
Gaia DR3 98692614681349248 2024.6869 0.070 0.055
Gaia DR3 990524191236435200 2024.8651 0.240 0.160
Gaia DR3 990524191236435456 2024.8651 0.195 0.135
Gaia DR3 992789150829982848 2024.1278 0.140 0.105
Gaia DR3 992789872384490496 2024.865 0.145 0.105